OpenStudy (anonymous):

f (x) = x 2 + 3x - 7 I need help

OpenStudy (nathan917):

Best Answer - Chosen by Voters 1. Assume that f(x) = (x/2) - 3x - 7 Substitute x with 5 to get: f(5) = 5/2 - 3(5) - 7 = 2.5 - 15 - 7 = 2.5 - 22 = -19.5 2. Do the same for this problem! f(5) = (5)² - 3(5) - 7 = 25 - 15 - 7 = 10 - 7 = 3

jhonyy9 (jhonyy9):

do you know formula for discriminat because this is a quadratic equation what i think that you need to solve it yes ?